OpenStudy (anonymous):

Taking this quickcheck right now.. Find all the zeros of each equation x3 – x2 + 5x – 5 = 0 a) isqrt5, –isqrt5, 1 b) isqrt5, –isqrt5, –1 c) sqrt5,-sqrt5, -1 d) isqrt5, 1 I suck at doing these. If you could show me the steps. I would appreiciate that.

OpenStudy (anonymous):

Start by factoring: You'll get something like (x-1)(x^2+5)=0 now solve each: (x-1)=0 (x^2+5)=0... solve this one using the quadratic formula.

OpenStudy (anonymous):

Start by solving the firt equation: (x-1)=0 x-1=0 x=1 now solve the second equation. (x^2+5)=0 x^2+5=0 subtract the 5 x^2=-5 take the square root of both sides....(no need to really use the quadratic formula). x=-sqrt5, x=-(isqrt5), x=1

OpenStudy (anonymous):

Oh ok I was confused on what the first step should be. So you factor first.. Ok. Thanks.